OpenStudy (anonymous):

Simplify (-3)^3 over (-3)^2. will give metal to best answer. A. -243 B. -3 C. 729 D. 9

OpenStudy (anonymous):

\[\frac{( -3)^3 }{ (-3)^2 }\]

jigglypuff314 (jigglypuff314):

equals (-3)^(3-2) = (-3)^1 = when it's division, the powers subtract

OpenStudy (anonymous):

umm, ok. so would it be B?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

thanks :)

jigglypuff314 (jigglypuff314):

you're welcome :)